### v1

@@ 3.3 Instruction Synthesis and Task Labelling
To enable instruction-driven modeling and facilitate stratified evaluation, each task in APE-Bench I is paired with a natural language instruction and annotated with semantic labels describing its functional category and structural difficulty. Both instruction synthesis and task labelling are performed automatically using LLMs, with verification procedures designed to ensure fidelity, clarity, and modeling relevance.
Instruction Generation.
For each filtered task derived from a Lean file diff (as described in Section 3.2), we synthesize an accompanying instruction that concisely captures the developer’s intent. We use Claude Sonnet 3.7 (thinking mode) to generate instructions from the raw diff, prompting the model to extract and summarize the intended edit in an imperative, concise, and actionable style. Instructions typically reflect changes such as introducing a new definition, modifying a proof, or restructuring an existing theorem.
To ensure semantic alignment between instruction and patch, we apply a reverse validation procedure identical to the semantic evaluation process described in Section 4.2. Specifically, the gold patch is re-evaluated against the synthesized instruction using the LLM-as-a-Judge framework. Only tasks for which the original edit satisfies the instruction are retained in the benchmark. This backward verification ensures that instructions are both meaningful and faithfully representative of the structural change.
Task Labelling.
Figure 4: Distribution of code change magnitudes across semantic labels within the test set. (a) Based on non-comment line counts from gold patches in the test set, grouped by difficulty. (b) Patch sizes grouped by functional category, based on gold commit diffs.
To enable stratified evaluation and reveal model capability differences across structurally diverse tasks, each instance in the test set is automatically annotated with semantic labels describing its functional intent and structural complexity. Specifically, we assign three types of labels:
  • Task Category, capturing functional purpose, is categorized as Feature, Refactor, or Bug Fix, aligned with common proof engineering operations.
  • Difficulty Level, estimating structural and reasoning complexity, is stratified as Very Easy, Easy, Medium, Hard, or Very Hard.
  • Task Nature, a binary label indicating whether a patch is Substantial or Superficial, used for filtering; only Substantial tasks are retained in the benchmark.
In addition to supporting dataset annotation, the labels also guide task selection: instances marked as Very Easy or Superficial are filtered out, ensuring that the benchmark retains only structurally substantive and semantically meaningful edits. Beyond this filtering role, the labels enable stratified evaluation across both functional categories and complexity levels, forming the foundation for the fine-grained performance analysis in Section 5.3. We next assess the effectiveness of these labels by analyzing their correlation with code magnitude and declaration-level structure.
Statistical Validation of Difficulty Labels.
Figure 4(a) shows that while Easy tasks are generally small, Medium and Hard tasks span a broad range of edit sizes, with no strict separation based on line count. This indicates that difficulty labels capture more than superficial metrics—they reflect underlying semantic and structural complexity. This is further confirmed by object-level patterns in Figure 5. Medium and Hard tasks involve more declarations and larger edits per object, particularly for theorems, suggesting greater logical coordination and proof effort. These trends align with empirical results in Section 5.3, where model performance declines sharply from Easy to Hard, validating the semantic significance of the difficulty labels.
Functional Differentiation by Task Category.
@@ 4 Evaluation Infrastructure and Protocol
Figure 6: Evaluation pipeline for APE-Bench I. A model-generated patch undergoes syntactic verification via the Lean compiler and semantic judgement via an LLM-as-a-Judge. A task is marked as successful only if both stages are passed.
To enable reliable and automated assessment of model-generated edits in APE-Bench I, we establish a two-stage evaluation framework that reflects the dual demands of proof engineering: syntactic correctness and semantic adequacy. Each candidate patch is evaluated for both grammatical validity under Lean’s type system and alignment with the task’s instruction-level intent.
The first stage, syntactic verification, checks whether the resulting file of the generated pass the Lean verification within the corresponding versioned environment. This process is implemented via Eleanstic, a scalable verification infrastructure that supports reproducible, parallelizable, and version-aware Lean verification infrastructure (Section 4.1).
The second stage, semantic judgement, determines whether the patch satisfies the instruction as intended. This is carried out by a large language model acting as an instruction-conditioned judge, which reasons over the pre-edit file, the patch, and the goal description to determine task success (Section 4.2).
Figure 6 presents the overall evaluation pipeline. A task is marked as successful only if a model-generated patch passes both verification stages. This protocol provides a standardized, interpretable, and scalable basis for computing success metrics, and enables detailed analysis of model performance across syntax-level and semantics-level axes.
@@ 4.1 Syntactic Verification via Eleanstic
Syntactic verification in APE-Bench I requires compiling model-generated patches within the exact versioned environment from which each task was derived. In large-scale formal libraries like Mathlib4, where definitions and dependencies evolve rapidly across commits, such version fidelity is critical to ensure meaningful evaluation. However, this fidelity introduces significant engineering complexity: naively rebuilding the library for every historical commit results in both prohibitive computational cost and unsustainable storage overhead. A single clean build of Mathlib4 can take several hours to complete and produce multiple gigabytes of intermediate artifacts. When scaled to thousands of commits—each potentially associated with thousands of patch candidates and evaluated across multiple models—this cost compounds rapidly, impeding both throughput and reproducibility.
To address this, we introduce Eleanstic, a dedicated infrastructure for scalable, version-aware syntactic verification. Eleanstic decouples the high-cost, stateful process of building Mathlib from the low-cost, stateless process of validating edits. It does so by collapsing each post-build state into a compact, content-addressable snapshot that can later be restored to yield a faithful Lean environment for verification. This architecture significantly reduces redundant computation, minimizes on-disk storage usage, and supports efficient parallel evaluation across thousands of distinct versions.
Snapshot Generation.
For each unique commit, Eleanstic first performs a clean lake build and recursively traverses the resulting worktree. Every file is hashed using SHA-256 and stored in a content-addressable storage (CAS) structure with a two-level hash-based directory layout. In parallel, a binary snapshot is created to record the mapping from relative file paths to content hashes and file types (regular or symbolic). This snapshot, encoded using fixed-width binary records, serves as a complete and portable blueprint of the post-build environment, ensuring reproducibility while minimizing metadata overhead.
Snapshot Restoration.
At verification time, the appropriate snapshot is loaded and its contents restored into a temporary, isolated worktree that exactly replicates the post-build file structure of the original compilation. Files are fetched from the global CAS and written to disk. This process is stateless and fully parallelizable, enabling fast switching across historical commits without requiring repeated builds. Empirically, Eleanstic restores a full Lean environment in under one second on SSD-backed systems. To support long-term scalability, Eleanstic performs aggressive global deduplication across all snapshots. Across 3734 commits spanning August 2023 to March 2025, it compresses 15.6 TB of build artifacts into just 1.1 TB, yielding a storage reduction of over 93%. This compression enables persistent caching of every historical state required for evaluation, without exhausting disk resources.
Patch Verification.
After restoring the environment, the model-generated patch (see Section 5.1) is applied to the target Lean file, and the resulting content is saved to a temporary file. Verification follows a strict file-level isolation policy: rooted in the restored Mathlib worktree, the Lean compiler compiles only the temporary file without affecting any files in the restored environment. Additionally, any temporary file that attempts to import the target file is discarded to prevent unintended leakage across file boundaries. A patch is considered syntactically valid if it compiles successfully within the corresponding version of Mathlib4 and does not emit warnings.
@@ 5.1 Experimental Setup
We evaluate seven general-purpose language models on APE-Bench I using a standardized evaluation pipeline. The task for each model is to generate a Lean patch that fulfills a natural language instruction, based on a complete pre-edit source file. Success is measured via a two-stage evaluation combining syntactic verification and semantic judgement.
Models.
We evaluate a range of state-of-the-art, general-purpose LLMs, selected to represent both long-context reasoning and standard generation capabilities:
  • Reasoning models: OpenAI o3-mini, Claude Sonnet 3.7 (thinking), DeepSeek R1.
  • Non-reasoning models: Claude Sonnet 3.7, DeepSeek V3 (250324), GPT-4o, Doubao 1.5 Pro.
All models are zero-shot, without any task-specific fine-tuning for proof engineering.
Patch Normalization via DiffRepair.
Model outputs are produced in unified diff format, which often contains structurally imprecise or misaligned edits (e.g., incorrect line anchors, out-of-date contexts). To ensure reliable application and verification, we introduce a post-processing phase called DiffRepair, which rewrites imprecise diffs into structurally consistent patches. This mechanism operates in two phases: first, it localizes the intended edit position using fuzzy anchor matching; then it reconstructs the patch with preserved structural and contextual integrity. All generated patches are normalized through DiffRepair before subsequent compilation or evaluation. (See Appendix A for full technical details.)
Sampling Methodology.
We adopt the standard pass@$k$ evaluation [4] to estimate task success rates. For each task, $n=20$ candidate patches are sampled, and pass@16 is computed using the unbiased estimator:
@@
where $c\leq n$ is the number of patches passing both syntactic and semantic verification.
Temperature Settings.
For models supporting temperature control, we evaluate outputs under two sampling temperatures, $T\in\{0.0,0.6\}$. Empirical results indicate that $T=0.6$ consistently improves output diversity and task coverage. For reasoning models that do not expose temperature parameters, we adopt default or recommended configurations: o3-mini is run with medium reasoning effort, while Claude Sonnet (thinking) is allocated 6,000 tokens of thinking budget to enable extended context comprehension.
@@ 5.2 Main Results
  • Lean Compile pass@16: The proportion of tasks for which at least one candidate compiles successfully using Lean 4, validating syntactic and structural soundness.
  • LLM-Judge pass@16: The proportion of tasks for which at least one compiled candidate additionally satisfies the instruction according to semantic evaluation via LLM-as-a-Judge.
  • Semantic Drop: The percentage decrease from Lean Compile to LLM-Judge success, reflecting how often syntactically valid outputs fail to meet the intended edits.
Table 2 summarizes the performance of seven general-purpose models. Among all candidates, o3-mini achieves the highest overall success rate, with 33.58% of tasks completed successfully under full evaluation. It also shows the highest syntactic pass rate at 48.38%, indicating its ability to generate compilable edits across a broad range of task types. However, its semantic drop-the proportion of compilable patches that fail to meet instruction-level requirements-is also substantial at over 30%, suggesting frequent deviation from intended modifications.
In contrast, Claude Sonnet 3.7 (thinking) exhibits perfect semantic consistency with a semantic drop of zero, yet achieves a significantly lower overall success rate due to reduced syntactic coverage. GPT-4o and Doubao 1.5 Pro show the opposite trend: both suffer from large semantic drops of 49.31% and 32.21%, respectively. These patterns highlight fundamental differences in model behavior—while some models excel at generating well-formed Lean code, they often fail to capture the semantic intent of the editing instruction.
Model Lean Compile pass@16 LLM-Judge pass@16 Semantic Drop (%)
o3-mini 48.38 33.58 30.58
Claude Sonnet 3.7 (thinking) 17.73 17.73 0.00
DeepSeek R1 17.67 15.85 10.30
DeepSeek V3 16.23 15.72 3.11
Claude Sonnet 3.7 15.85 14.85 6.31
GPT-4o 18.94 9.60 49.31
Doubao 1.5 Pro 9.70 6.58 32.21
Table 2: Overall model performance on APE-Bench I using pass@16. Semantic Drop is computed as the percentage of Lean-compilable outputs that fail LLM-based semantic validation.
Figure 7 further illustrates model behavior across different pass@$k$ thresholds. The left panel reports syntactic success (Lean Compile), while the right panel shows semantic success (LLM-Judge) conditioned on compilation. o3-mini consistently outperforms other models across $k$ values in both dimensions, confirming its broad sampling diversity and task coverage. Claude Sonnet 3.7 (thinking) remains flat across $k$, reflecting its low-variance generation strategy.
Figure 7: Model performance across pass@$k$ values. Left: Lean Compile pass@$k$ (syntactic validity); Right: LLM Judge pass@$k$ (semantic correctness conditioned on compilation). The gap between the two reflects semantic degradation.
Notably, even the best-performing model—o3-mini—succeeds on only about one-third of the benchmark tasks, despite the fact that all tasks are constrained to single-file edits of no more than 100 lines. This highlights the substantial gap that remains between present capabilities and the demands of real-world proof engineering. In the following sections, we analyze this failure spectrum in greater detail by examining the structural and semantic robustness of model outputs.
@@ 5.3 Structural Robustness and Task-Scale Sensitivity
While overall success rates reveal substantial variation across models, they do not capture how model performance scales with the structural complexity of the underlying tasks. In this section, we analyze the extent to which models can handle edits of increasing size and structural scope, assess their ability to produce appropriately scaled patches, and evaluate their robustness across task types, difficulty levels, and mathematical domains.
Performance Degradation on Larger Edits.

### v2

@@ 3.3 Instruction Synthesis and Task Labelling
Figure 4: Distribution of code change magnitudes across semantic labels in the test set. (a) Based on non-comment line counts from gold patches, grouped by difficulty. (b) Patch sizes grouped by functional category.
Each task is paired with a natural language instruction and annotated with semantic labels describing its functional category and structural difficulty. For instruction generation, we use Claude Sonnet 3.7 (thinking mode) to extract and summarise the developer’s intent from the raw diff in an imperative style. We apply a reverse validation procedure to ensure semantic alignment between instruction and patch (detailed methodology in Appendix B.2).
For task labelling, we assign three types of labels to the test set:
  • Task Category: Feature, Refactor, or Bug Fix, aligned with common proof engineering operations.
  • Difficulty Level: Very Easy, Easy, Medium, Hard, or Very Hard, reflecting increasing structural and logical complexity. Very Easy tasks are filtered out.
  • Task Nature: Substantial or Superficial, used for filtering; only Substantial tasks are retained.
Figure 4 shows the distribution of code changes across difficulty levels and functional categories. These labels enable stratified evaluation and help reveal capability gaps across models, as confirmed by our experimental results in Section 5. Appendix B.3 provides additional statistical validation of these label distributions.
@@ 4 Evaluation Infrastructure and Protocol
To enable reliable assessment of model-generated edits in APE-Bench I, we establish a two-stage evaluation framework that reflects the dual demands of proof engineering: syntactic correctness and semantic adequacy. As illustrated in Figure 5, each candidate patch is evaluated for both grammatical validity under Lean’s type system and alignment with the task’s instruction-level intent. The first stage, syntactic verification, checks whether the resulting file passes Lean verification within the corresponding versioned environment. This process is implemented via Eleanstic, a scalable verification infrastructure that supports reproducible, parallelisable, and version-aware Lean verification (Section 4.1). The second stage, semantic judgement, determines whether the patch satisfies the instruction as intended. This is carried out by a large language model acting as an instruction-conditioned judge (Section 4.2). A task is marked as successful only if a model-generated patch passes both verification stages. This protocol provides a standardised and scalable basis for computing success metrics, enabling detailed analyses across syntax-level and semantics-level dimensions.
@@ 4.1 Syntactic Verification via Eleanstic
Figure 5: Evaluation pipeline for APE-Bench I. Given a commit hash, the corresponding Lean environment is restored by a snapshot from content-addressable storage (CAS), reproducing the exact worktree state. An LLM generates a patch based on the instruction and the original file. This patch is applied to a temporary file and evaluated through Lean compile and LLM judge.
Syntactic verification in APE-Bench I requires compiling model-generated patches within the exact versioned environment from which each task was derived. In large-scale formal libraries like Mathlib4, where definitions and dependencies evolve rapidly across commits, such version fidelity is critical for meaningful evaluation.
We introduce Eleanstic, a dedicated infrastructure for scalable, version-aware syntactic verification. Eleanstic decouples the high-cost process of building Mathlib4 from the low-cost process of validating edits by collapsing each post-build state into a compact, content-addressable snapshot. This architecture significantly reduces redundant computation, minimises storage usage, and supports efficient parallel evaluation across thousands of distinct versions (technical implementation details in Appendix B.4).
For verification, a model-generated patch is applied to the target Lean file in an isolated environment, and the Lean compiler verifies only the modified file without affecting the restored environment. A patch is considered syntactically valid if it compiles successfully and does not emit warnings.
@@ 5.1 Experimental Setup
We evaluate eight state-of-the-art language models on APE-Bench I: Gemini 2.5 Pro Preview, o3-mini, Claude Sonnet 3.7 (with and without thinking mode), DeepSeek R1, DeepSeek V3, GPT-4o, and Doubao 1.5 Pro. All models generate patches in unified diff format without task-specific fine-tuning. We process outputs using DiffRepair, a post-processing mechanism that rewrites imprecise diffs into structurally consistent patches (described in Appendix C.3). We adopt the standard pass@$k$ evaluation [3] to estimate task success rates. For each task, $n=20$ candidate patches are sampled, and pass@16 is computed using the unbiased estimator:
@@
where $c\leq n$ is the number of patches passing both syntactic and semantic judgement. For most models, we use temperature T=0.6 to improve output diversity.
@@ 5.2 Main Results
  • Verification pass@16: The proportion of tasks for which at least one candidate compiles successfully using Lean 4.
  • Judgement pass@16: The proportion of tasks for which at least one compiled candidate additionally satisfies the instruction according to semantic evaluation.
  • Relative Decrease: The percentage decrease from Verification to Judgement success, reflecting how often syntactically valid outputs fail to meet the intended edits.
Model Verification pass@16 Judgement pass@16 Relative Decrease (%)
Gemini 2.5 Pro Preview 18.79% 18.04% 3.99%
Claude Sonnet 3.7 (thinking) 12.73% 12.73% 0.001%
DeepSeek R1 15.30% 12.55% 17.97%
DeepSeek V3 12.44% 11.81% 5.06%
Claude Sonnet 3.7 11.33% 10.83% 4.41%
o3-mini 20.13% 8.60% 57.28%
GPT-4o 15.16% 6.73% 55.61%
Doubao 1.5 Pro 6.19% 2.98% 51.86%
Table 2: Overall model performance on APE-Bench I using pass@16. Relative Decrease is computed as the percentage of Lean-compilable outputs that fail LLM-based semantic judgement.
As shown in Table 2, Gemini 2.5 Pro Preview achieves the highest judgement success rate at 18.04%, while o3-mini leads in verification pass rate (20.13%) but suffers from a substantial relative decrease (57.28%). Claude Sonnet 3.7 (thinking) exhibits nearly perfect semantic consistency with a negligible relative decrease (0.001%), though at a lower overall success rate. Notably, even the best-performing model succeeds on only about one-sixth of the benchmark tasks, highlighting the substantial gap between present capabilities and the demands of real-world proof engineering. Additional performance metrics across different pass@k values are available in C.
@@ 5.3 Structural Robustness and Task-Scale Sensitivity
Figure 6: Model pass@16 success rate grouped by ground-truth patch size. Performance declines rapidly with increasing patch size. Gemini 2.5 Pro Preview leads in small to medium ranges, but no model demonstrates robustness beyond 50-line edits.